OpenStudy (gavin39):

Identify whether the series summation of 15 open parentheses 4 close parentheses to the I minus 1 power from 1 to infinity is a convergent or divergent geometric series and find the sum, if possible. This is a convergent geometric series. The sum is –5. This is a divergent geometric series. The sum is –5. This is a convergent geometric series. The sum cannot be found. This is a divergent geometric series. The sum cannot be found.

OpenStudy (gavin39):

I believe its C

OpenStudy (anonymous):

still trying to read it an equation editor might help

OpenStudy (anonymous):

be that as it may, if it is four to some power, it is geometric, but since \(4\geq 1\) it is not summable

OpenStudy (gavin39):

for some reason not letting me put it in

OpenStudy (kirbykirby):

Do you mean something like this: \[\sum_{i=1}^{\infty}15(4)^{i-1}\]?

OpenStudy (gavin39):

yes!

OpenStudy (kirbykirby):

@satellite73 is right though, it's a geometric series with \(|r|=4\ge 1\) which won't converge!

OpenStudy (anonymous):

so do you know what the right answer was? i hate this kinda math its horrible!

OpenStudy (kirbykirby):

If it's divergent, then you can't find its sum

OpenStudy (anonymous):

Thank you sooo much!!!